OpenStudy (anonymous):

Help me please (x^2 + 2x)/(12x + 54) - (3 - x)/(8x +36)

OpenStudy (compassionate):

Hello, your first step is to factor the expression. Can you do that for me, hun?

OpenStudy (anonymous):

wait first aren't you supposed to turn the subtraction sign into an addition sign... and you do that by multiplying it by -1 and it only effects the numerator

OpenStudy (compassionate):

No - your first step is simplifying.

OpenStudy (compassionate):

I actually have to go! Sorry! But I bet @ganeshie8 or @goformit100 can assist you! :)

OpenStudy (anonymous):

x(x + 2)/ 6(2x + 9) - (-3 + x)/4(2x + 9)?

ganeshie8 (ganeshie8):

you mean x(x + 2)/ 6(2x + 9) + (-3 + x)/4(2x + 9)?

ganeshie8 (ganeshie8):

\(\large \frac{x(x + 2)}{ 6(2x + 9)} + \frac{(-3 + x)}{4(2x + 9)} \)

OpenStudy (anonymous):

oops ya sorry.. and now I know we're supposed to have a common denominator

ganeshie8 (ganeshie8):

yes, so how to get a common denominator ?

ganeshie8 (ganeshie8):

wats lcm of 6 and 4 ?

ganeshie8 (ganeshie8):

is it 12 ?

OpenStudy (anonymous):

You multiply 4 to the numerator and the denominator of the first fraction and multiply 6 to the numerator and the denominator of the 2nd fraction?

ganeshie8 (ganeshie8):

perfect !

ganeshie8 (ganeshie8):

\(\large \frac{x(x + 2)}{ 6(2x + 9)} + \frac{(-3 + x)}{4(2x + 9)} \) \(\large \frac{4x(x + 2)}{ 24(2x + 9)} + \frac{6(-3 + x)}{24(2x + 9)} \)

ganeshie8 (ganeshie8):

now that we have denominators same, simply add numerators

OpenStudy (anonymous):

ok that makes sense so far

OpenStudy (anonymous):

would you multiply the 6 and the 4x in the numerators

ganeshie8 (ganeshie8):

yes

ganeshie8 (ganeshie8):

\(\large \frac{x(x + 2)}{ 6(2x + 9)} + \frac{(-3 + x)}{4(2x + 9)} \) \(\large \frac{4x(x + 2)}{ 24(2x + 9)} + \frac{6(-3 + x)}{24(2x + 9)} \) \(\large \frac{4x^2+8x}{ 24(2x + 9)} + \frac{-18 + 6x}{24(2x + 9)} \)

ganeshie8 (ganeshie8):

add them now

ganeshie8 (ganeshie8):

\(\large \frac{x(x + 2)}{ 6(2x + 9)} + \frac{(-3 + x)}{4(2x + 9)} \) \(\large \frac{4x(x + 2)}{ 24(2x + 9)} + \frac{6(-3 + x)}{24(2x + 9)} \) \(\large \frac{4x^2+8x}{ 24(2x + 9)} + \frac{-18 + 6x}{24(2x + 9)} \) \(\large \frac{4x^2+8x-18+6x}{ 24(2x + 9)} \)

ganeshie8 (ganeshie8):

simplify

OpenStudy (anonymous):

why wouldn't you just multiply the 6 and 4x and just leave it like 24x(x + 2)(x - 3)

OpenStudy (anonymous):

could you do that?

ganeshie8 (ganeshie8):

dint get u

OpenStudy (anonymous):

for the 4x and 6 you distributed but do you have to do that or can you just combine them like 24x(x +2)(-3 + x)

OpenStudy (anonymous):

?

ganeshie8 (ganeshie8):

you will have to distribute that

ganeshie8 (ganeshie8):

\(\large \frac{x(x + 2)}{ 6(2x + 9)} + \frac{(-3 + x)}{4(2x + 9)} \) \(\large \frac{4x(x + 2)}{ 24(2x + 9)} + \frac{6(-3 + x)}{24(2x + 9)} \) \(\large \frac{4x^2+8x}{ 24(2x + 9)} + \frac{-18 + 6x}{24(2x + 9)} \) \(\large \frac{4x^2+8x-18+6x}{ 24(2x + 9)} \) \(\large \frac{4x^2+14x-18}{ 24(2x + 9)} \) \(\large \frac{2(2x^2+7x-9)}{ 24(2x + 9)} \) \(\large \frac{2x^2+7x-9}{ 12(2x + 9)} \)

ganeshie8 (ganeshie8):

we end up like that see if it makes some sense

OpenStudy (anonymous):

okay so we HAVE to distribute

ganeshie8 (ganeshie8):

yup !

ganeshie8 (ganeshie8):

see if u can factor numerator and can cancel something

OpenStudy (anonymous):

(2x^2 - 2x)(9x - 9)

ganeshie8 (ganeshie8):

\(\large \frac{x(x + 2)}{ 6(2x + 9)} + \frac{(-3 + x)}{4(2x + 9)} \) \(\large \frac{4x(x + 2)}{ 24(2x + 9)} + \frac{6(-3 + x)}{24(2x + 9)} \) \(\large \frac{4x^2+8x}{ 24(2x + 9)} + \frac{-18 + 6x}{24(2x + 9)} \) \(\large \frac{4x^2+8x-18+6x}{ 24(2x + 9)} \) \(\large \frac{4x^2+14x-18}{ 24(2x + 9)} \) \(\large \frac{2(2x^2+7x-9)}{ 24(2x + 9)} \) \(\large \frac{2x^2+7x-9}{ 12(2x + 9)} \) \(\large \frac{(x-1)(2x+9)}{ 12(2x + 9)} \) \(\large \frac{x-1}{ 12} \)

OpenStudy (anonymous):

that makes soo much sense thank you =)

ganeshie8 (ganeshie8):

np :)

OpenStudy (anonymous):

wait